### v1

@@ Abstract
Recent progress in large language models (LLMs) has shown promise in formal theorem proving, yet existing benchmarks remain limited to isolated, static proof tasks, failing to capture the iterative, engineering-intensive workflows of real-world formal mathematics libraries. Motivated by analogous advances in software engineering, we introduce the paradigm of Automated Proof Engineering (APE), which aims to automate proof engineering tasks such as feature addition, proof refactoring, and bug fixing using LLMs. To facilitate research in this direction, we present APE-Bench I, the first realistic benchmark built from real-world commit histories of Mathlib4, featuring diverse file-level tasks described in natural language and verified via a hybrid approach combining the Lean compiler and LLM-as-a-Judge. We further develop Eleanstic, a scalable parallel verification infrastructure optimized for proof checking across multiple versions of Mathlib. Empirical results on state-of-the-art LLMs demonstrate strong performance on localized edits but substantial degradation on handling complex proof engineering. This work lays the foundation for developing agentic workflows in proof engineering, with future benchmarks targeting multi-file coordination, project-scale verification, and autonomous agents capable of planning, editing, and repairing formal libraries.
@@ 1 Introduction
Large language models (LLMs) have recently demonstrated impressive capabilities in formal mathematics, achieving state-of-the-art results on isolated theorem-proving tasks. Benchmarks such as miniF2F [17] and competition-grade problems like those in the IMO [5] showcase LLMs’ ability to reason symbolically and generate formally valid proofs. However, these benchmarks operate in static, goal-driven settings and fail to reflect the full complexity of real-world formal mathematics workflows. In practice, maintaining large formal libraries such as Mathlib4 entails more than solving standalone problems—it involves iteratively refactoring, extending, and debugging a highly interconnected and evolving codebase. This gap between isolated problem-solving and real-world proof engineering highlights the need for a fundamentally different evaluation paradigm.
@@
While analogous limitations in code generation have led to increasingly realistic software engineering benchmarks—e.g., from HumanEval [4] to SWE-bench [10]—the formal mathematics community has yet to adopt a similar shift. In software engineering, SWE-bench reframes model evaluation as resolving realistic GitHub issues grounded in historical code evolution. We argue that formal mathematics must undergo a comparable transformation: from static theorem proving to full-fledged proof engineering, where modifications are driven by long-term development and maintenance goals rather than isolated logical challenges. Contributors to large formal libraries such as Mathlib4 engage in continuous restructuring, extension, and repair across an evolving codebase. To model this process in a way that supports LLM-based evaluation, we introduce the paradigm of Automated Proof Engineering (APE)—a setting where each task presents a natural language instruction along with a pre-edit version of Lean files, and the model is expected to generate a corresponding code patch. While human edits are not guided by explicit instructions, we distill their underlying intent into instruction–patch pairs, enabling modular, interpretable, and scalable evaluation units. Indeed, the name APE intentionally alludes to this mimicry: we aim to “ape” the structure of human proof engineering behavior in a form suitable for language models.
@@
To operationalize this paradigm, we introduce the APE-Bench series—a staged suite of benchmarks designed to evaluate increasingly sophisticated proof engineering capabilities. As shown in Figure 1, this series progresses through three stages. APE-Bench I focuses on single-file instruction-guided edits with localized structural scope. APE-Bench II will introduce multi-file coordination and project-level verification, enabling evaluation of consistency-aware editing across interdependent modules. Finally, APE-Bench III aims to support autonomous workflows, where agents plan, edit, and iteratively repair formal libraries via feedback-driven strategies. This roadmap is not merely a scaling of difficulty, but a reflection of the layered competencies required for realistic formal proof maintenance—from local patch generation to cross-file planning and agentic self-correction.
@@
Figure 1: The APE-Bench series: A staged roadmap for automated proof engineering. The benchmark suite progresses from single-file edits (I) to multi-file coordination (II) and ultimately to autonomous, feedback-driven workflows (III). This staged design reflects the increasing structural and semantic complexity of real-world proof engineering tasks. Dashed arrows denote prospective support for feedback-driven self-correction.
Figure 2: APE-Bench I: Benchmark structure and evaluation pipeline. (a) Unlike earlier benchmarks such as miniF2F and miniCTX, which focus on isolated theorem-proving in minimal contexts, APE-Bench I targets realistic file-level proof engineering tasks sourced from real-world Mathlib commits. These tasks involve rich, interleaved Lean declarations and demand long-range reasoning and structural precision. (b) The data pipeline begins with commit mining, diff extraction, and instruction synthesis, followed by a one-time build phase where each Mathlib version is compiled and stored using Eleanstic—a multi-version, content-addressable infrastructure for efficient Lean verification. Each commit is snapshotted as a mapping from file paths to content hashes, with all unique content stored in a global deduplicated store. During evaluation, model-generated patches are applied in isolated worktrees and verified via a two-stage process: syntactic checking with the Lean compiler and semantic checking using an LLM-as-a-Judge. This setup supports scalable, parallel, and version-aware evaluation of proof engineering capabilities.
@@
This paper presents APE-Bench I, the first step in this roadmap. Built from thousands of real-world commits in Mathlib4, it targets localized, instruction-driven edits at the file level, providing a realistic and controlled setting for evaluating proof engineering capabilities. Its design is guided by four core goals:
  • Realism: Reflect real development practices in large-scale formal libraries;
  • Stratification: Support diverse task types and structural complexities to reveal capability gaps and preserve headroom;
  • Scalable Syntax-Semantics Evaluation: Enable syntactically efficient and semantically faithful assessment of candidate edits;
@@
  • Continual Evolvability: Allow automated updates in step with the evolution of the underlying library.
To realize realism, we extract file-level edits from real commits in Mathlib4, preserving their structural and contextual integrity. Each edit is paired with a natural language instruction synthesized from the diff and validated through a reverse LLM-based semantic check to ensure fidelity to the original developer intent.
To ensure stratification, we construct a high-quality and diverse task set through a multi-stage filtering pipeline that removes trivial or structurally incoherent edits. To expose and verify the structural diversity of the dataset, we automatically label each task by both functional category (e.g., feature, refactor, bug fix) and difficulty level, using LLMs. These labels are empirically validated: model performance across types and difficulty bands confirms their effectiveness in revealing capability differences and guiding fine-grained evaluation.
To support scalable and semantically grounded evaluation, we adopt a two-stage protocol: syntactic checking via the Lean compiler in isolated environments (enabled by our Eleanstic infrastructure), and semantic verification via an LLM-as-a-Judge system assessing whether the patch fulfills the instruction. This approach balances robustness, automation, and evaluation cost.
Finally, to ensure continual evolvability, the entire benchmark construction pipeline—from commit mining to task synthesis, labeling, and verification—is fully automated and publicly released. It can be re-executed over future Mathlib4 versions, enabling the benchmark to remain current without manual intervention.
APE-Bench I enables a comprehensive evaluation of leading language models on realistic proof engineering tasks, revealing clear performance gaps and structural limitations. o3-mini leads by a wide margin, achieving nearly twice the semantic success rate of the next best model. While most models degrade sharply as task size increases, o3-mini remains effective even on large edits. Many models frequently produce syntactically valid but semantically incorrect patches, failing to satisfy the intended instruction. In contrast, Claude Sonnet 3.7 (thinking) achieves high semantic precision but with limited coverage. To guide future research, we categorize common semantic failure types and provide case studies that highlight recurring issues in structural planning and task interpretation.
@@
    We introduce the paradigm of Automated Proof Engineering (APE), which reframes formal mathematics as instruction-driven code editing aligned with real-world development and maintenance workflows.
@@
    We construct APE-Bench I, the first large-scale, file-level benchmark for proof engineering, built from over 10,000 real-world commits in Mathlib4. Tasks are automatically synthesized by mining commit diffs and generating natural-language instructions, then labeled and validated to reflect realistic intent and structural difficulty.
@@
    We develop a two-stage evaluation protocol that combines syntax-level checking via Lean compilation with semantic-level verification using LLMs. To support this, we introduce Eleanstic, a scalable, version-aware Lean infrastructure that enables efficient, stateless patch validation across thousands of commits.
@@
    We conduct a comprehensive empirical study on leading LLMs, revealing substantial performance differences, sharp degradation on structurally complex tasks, and frequent semantic failures. We further provide a taxonomy of error types and case studies to guide future research.
@@
    We will release all data, code, and infrastructure to support reproducible evaluation and continual benchmark expansion as Mathlib4 evolves.
@@ 2 Background and Related Work
2.1 Formal Proof Engineering: Practice and Challenges
Formal proof engineering refers to the disciplined application of software engineering principles to the construction, maintenance, and evolution of large-scale formal mathematical libraries [14]. In contrast to traditional theorem proving, which typically centers on isolated logical goals, proof engineering entails a broader and more structurally intricate set of tasks—such as extending libraries with new features, restructuring existing proofs (refactoring), and correcting erroneous logic (bug fixing).
Real-world formal systems, including Lean’s Mathlib [13], Isabelle/HOL’s Archive of Formal Proofs, and the seL4 microkernel verification project [11], exemplify the scale and complexity of this engineering effort. These libraries often contain thousands of interdependent definitions, theorems, and proofs, where modifying a single element may trigger cascading changes across the codebase. Ensuring logical consistency under such conditions demands significant manual labor and deep structural awareness.
While certain tools—such as iCoq for regression-proving [3]—provide partial automation, they are largely heuristic and require intensive human oversight. To date, there remains a lack of automated systems capable of performing high-fidelity edits in complex proof environments.
2.2 Machine Learning for Theorem Proving
Early machine learning efforts in formal reasoning primarily focused on tactic prediction, where proof construction is modeled as a sequential decision-making process. Systems such as HOList [2] and GamePad [8] trained neural models to select proof tactics from localized proof contexts. However, these benchmarks rely on tasks extracted from existing libraries, often involving subgoals that lack standalone mathematical significance. Consequently, they offer only limited insight into a model’s genuine reasoning capabilities.
Recent research has shifted toward end-to-end theorem proving, where models are evaluated on their ability to solve self-contained problems without intermediate scaffolding. Datasets like miniF2F [17] and FIMO [12] reflect this transition and represent a meaningful step toward assessing mathematical competence.
2.3 Limitations of Existing Benchmarks
Nevertheless, most existing benchmarks still emphasize single-goal theorem proving within static and minimal contexts, a setting that diverges significantly from the demands of real-world proof maintenance. These benchmarks evaluate success based on isolated proof completion, neglecting the broader engineering workflows needed to evolve and maintain large formal mathematics libraries.
Table 1 summarizes the conceptual differences between conventional theorem-proving benchmarks and practical proof engineering settings. In contrast to the narrow and static focus of existing benchmarks, real-world workflows involve continuous updates, structural consistency, and diverse goals such as refactoring, bug fixing, and extending functionality.
@@ 2.3 Limitations of Existing Benchmarks
Objectives Single-attempt theorem completion Iterative refinement, multiple engineering goals
Activities Individual proof completion Feature addition, proof refactoring, bug fixing
Table 1: Comparison between conventional theorem-proving benchmarks and real-world proof engineering practices.
Several recent efforts have explored richer proof generation scenarios, such as incorporating intermediate steps, leveraging auxiliary definitions [6, 9, 15, 16, 7], or modifying existing proofs through refactoring [18] and improvement techniques [1]. However, these approaches are still typically used as tools for data augmentation or pipeline enhancement, rather than serving as rigorous evaluation frameworks. As a result, a clear gap remains in the form of systematic, task-diverse benchmarks that reflect the full scope of real-world proof engineering.
@@ 3 APE-Bench I: A Realistic Benchmark for File-Level Proof Engineering
To concretely realize the paradigm of Automated Proof Engineering (APE), we introduce APE-Bench I, a realistic and instruction-driven benchmark constructed from real-world edits in the Mathlib4 formal mathematics library. This section describes the design and construction of the benchmark, including its task format, the methodology for extracting and filtering file-level edits from version history, and the automatic synthesis of natural-language instructions and semantic labels. Together, these components enable structured, modular, and semantically faithful evaluation of language models on localized proof engineering tasks.
@@ 3.1 Task Format and Objective
Each task in APE-Bench I defines a localized file-level proof engineering problem, grounded in an actual edit to a Lean source file. The objective is to apply a valid patch that fulfills a natural-language instruction by transforming a given pre-edit file into a structurally correct and semantically meaningful post-edit version.
@@
  • Patch is a unified diff that encodes the intended edit, specifying insertions and deletions relative to PreFile. The diff format supports modular and interpretable editing, and facilitates automated application, verification, and evaluation.
This triplet formulation reflects the real-world dynamics of proof maintenance, where developers operate on full files with evolving structures rather than isolated subgoals. It also aligns naturally with instruction-following language models, enabling instruction-conditioned generation and evaluation of file-level structural changes. By grounding each task in a concrete diff from version history, APE-Bench I preserves the realism, coherence, and modularity essential for evaluating automated proof engineering capabilities.
3.2 Task Extraction from Mathlib Commits
(a) Token distribution by benchmark
@@ 3.2 Task Extraction from Mathlib Commits
Figure 3: APE-Bench I: Scale and complexity of extracted tasks. (a) Compared to prior datasets, APE-Bench I tasks span significantly longer input–output contexts, reflecting greater structural and semantic complexity. (b) Starting from 19k commits, the extraction pipeline produces over 185k candidates, which are refined through multiple filtering stages to yield 10,928 high-quality tasks.
To construct realistic and structurally diverse proof engineering tasks, we extract file-level edits from the version history of Mathlib4, Lean’s primary formal mathematics library. The resulting benchmark tasks are derived from actual developer activity, ensuring contextual fidelity and development relevance. This section describes the extraction pipeline, which proceeds in four stages: commit sourcing and decomposition, chunk-level expansion, multi-stage filtering, and syntactic validation.
@@
We begin with 19,867 Git commits made between August 2023 and March 2025, a period selected to capture stable post-migration development in Mathlib4. Each commit is decomposed into file-level diffs, isolating modified files and their corresponding changes. These diffs are further segmented into semantically scoped chunks: for modified files, we use hunk-level boundaries; for newly added files, we apply heuristics based on Lean declaration structure to determine logical segments.
@@
To improve modeling tractability and enhance semantic interpretability, we decompose large diffs into smaller, contiguous edit fragments. Each fragment is applied incrementally to the evolving file state, yielding fine-grained and contextually grounded task units. This expansion increases the total number of edit candidates by nearly an order of magnitude, producing over 187,000 candidate units from the original commit set.
@@
To isolate high-quality, semantically meaningful tasks, we apply a sequence of automated filters targeting three dimensions:
  • Semantic relevance: We retain only commits with recognizable editing intent, identified via conventional prefixes such as feat, fix, or refactor. Commits involving file deletions, mechanical formatting, or edits outside the Mathlib/ directory are excluded.
  • Content quality: Edits that only affect comments, import ordering, or whitespace are discarded, as they lack structural substance.
  • Structural integrity: We restrict each task to at most 100 lines of non-comment changes and a maximum input length of 16,000 tokens (measured using the GPT-4o tokenizer). Tasks must be localized, coherent, and semantically self-contained.
The most substantial reduction occurs after chunking, where structurally trivial diffs are eliminated. Subsequent filters further refine the corpus. The overall retention trajectory is visualized in Figure 3(b), illustrating each pruning stage in the pipeline.
@@
To ensure structural soundness, all retained patches are compiled under their post-edit environments using Lean. Only those patches that successfully typecheck in the restored version context are included in the final benchmark. This step guarantees that each task is grounded in a valid and verifiable edit to a Lean file.
The final benchmark consists of 10,928 structurally coherent file-level tasks. As shown in Figure 3(a), these tasks exhibit significantly longer input and output sequences than prior benchmarks such as miniF2F and miniCTX. This reflects the higher structural complexity of proof engineering tasks and highlights the need for long-context reasoning and precise structural planning.
Each stage of the pipeline incrementally refines the dataset to ensure structural integrity, semantic depth, and fidelity to real-world development practices. The resulting benchmark comprises 10,928 structurally coherent file-level tasks that capture a diverse set of realistic, non-trivial modifications authored by human contributors. This dataset serves as the foundation for instruction synthesis and stratified evaluation in the subsequent sections. As shown in Figure 3(a), these tasks exhibit significantly longer input and output sequences than prior benchmarks such as miniF2F and miniCTX, underscoring the greater structural complexity of proof engineering and the need for long-context reasoning and precise structural planning.

### v2

@@ Abstract
Recent progress in large language models (LLMs) has shown promise in formal theorem proving, yet existing benchmarks remain limited to isolated, static proof tasks, failing to capture the iterative, engineering-intensive workflows of real-world formal mathematics libraries. Motivated by analogous advances in software engineering, we introduce the paradigm of Automated Proof Engineering (APE), which aims to automate proof engineering tasks such as feature addition, proof refactoring, and bug fixing using LLMs. To facilitate research in this direction, we present APE-Bench I, the first realistic benchmark built from real-world commit histories of Mathlib4, featuring diverse file-level tasks described in natural language and verified via a hybrid approach combining the Lean compiler and LLM-as-a-Judge. We further develop Eleanstic, a scalable parallel verification infrastructure optimized for proof checking across multiple versions of Mathlib. Empirical results on state-of-the-art LLMs reveal a syntax–semantics gap: while some models perform well on localized edits, overall success rates drop sharply on structurally complex tasks. This work lays the foundation for developing agentic workflows in proof engineering, with future benchmarks targeting multi-file coordination, project-scale verification, and autonomous agents capable of planning, editing, and repairing formal libraries.
@@ 1 Introduction
Large language models (LLMs) have recently demonstrated impressive capabilities in formal mathematics, achieving state-of-the-art results on isolated theorem-proving tasks. However, existing benchmarks operate in static, goal-driven settings and fail to reflect the full complexity of real-world formal mathematics workflows. Maintaining large formal libraries such as Mathlib4 entails more than solving standalone problems—it involves iteratively refactoring, extending, and debugging an interconnected and evolving codebase. This gap between isolated problem-solving and real-world proof engineering highlights the need for a fundamentally different evaluation paradigm.
@@
While analogous limitations in code generation have led to increasingly realistic software engineering benchmarks-e.g., from HumanEval [3] to SWE-bench [7]-the formal mathematics community has yet to adopt a similar shift. We introduce the paradigm of Automated Proof Engineering (APE)—a setting where each task presents a natural language instruction along with a pre-edit version of Lean files, and the model generates a corresponding code patch. While human edits are not guided by explicit instructions, we distill their underlying intent into instruction–patch pairs, enabling interpretable and scalable evaluation units.
@@
To operationalise this paradigm, we introduce the APE-Bench series—a staged suite of benchmarks for increasingly sophisticated proof engineering capabilities. As shown in Figure 1, APE-Bench I focuses on single-file instruction-guided edits with localised scope. APE-Bench II will introduce multi-file coordination and project-level verification. Finally, APE-Bench III aims to support autonomous workflows, where agents plan, edit, and iteratively repair formal libraries via feedback-driven strategies.
@@
Figure 1: The APE-Bench series: A staged roadmap for automated proof engineering. The benchmark suite progresses from single-file edits (APE-Bench I) to multi-file coordination (APE-Bench II) and ultimately to autonomous workflows (APE-Bench III). This staged design reflects the increasing complexity of real-world proof engineering tasks.
Figure 2: APE-Bench I: Benchmark structure and evaluation pipeline. (a) Unlike earlier benchmarks focusing on isolated theorem-proving, APE-Bench I targets realistic file-level proof engineering tasks from real-world Mathlib commits. (b) The data pipeline includes commit mining, diff extraction, instruction synthesis, and verification using Eleanstic-a multi-version infrastructure for efficient Lean verification.
@@
This paper presents APE-Bench I, the first step in this roadmap. Built from thousands of real-world commits in Mathlib4, it targets localised, instruction-driven edits at the file level. Its design is guided by four core goals:
  • Realism: Reflect real development practices by extracting file-level edits from actual Mathlib4 commits, preserving their structural integrity.
  • Stratification: Support diverse task types and complexities through a multi-stage filtering pipeline that removes trivial edits and categorizes tasks by function and difficulty.
  • Scalable Syntax-Semantics Evaluation: Enable efficient assessment through syntactic verification via the Lean compiler and semantic judgement via an LLM-as-a-judge system.
@@
  • Continual Evolvability: Support automated updates with the evolution of the underlying library through fully automated construction pipelines.
APE-Bench I provides a rigorous framework for evaluating state-of-the-art language models on realistic proof engineering tasks. Experimental results show that although Gemini 2.5 Pro Preview achieves the highest semantic accuracy under LLM-based evaluation, it successfully completes only one-sixth of the benchmark tasks and fails entirely on the most challenging ones. Most models exhibit significant performance degradation as task complexity increases, often generating patches that are syntactically valid but semantically flawed. These findings underscore a substantial gap between current model capabilities and the demands of real-world proof engineering.
@@
    We present Automated Proof Engineering (APE)—a paradigm that casts LLM-based formal reasoning as instruction-driven code edits reflecting real-world development practices.
@@
    We construct APE-Bench I, the first large-scale, file-level benchmark for proof engineering, built from over 10,000 real-world commits in Mathlib4.
@@
    We develop a two-stage evaluation protocol combining syntactic verification via Lean compilation with semantic judgement via an LLM-as-a-judge system.
@@
    We conduct a comprehensive empirical study on state-of-the-art LLMs, revealing substantial performance differences, sharp degradation on complex tasks, and frequent semantic failures.
@@
    We release all data, code, and infrastructure to support reproducible evaluation and continual benchmark expansion as Mathlib4 evolves.
@@ 2 Background and Related Work
Formal proof engineering refers to the disciplined application of software engineering principles to the construction, maintenance, and evolution of large-scale formal mathematical libraries [10]. In contrast to traditional theorem proving, which typically centres on isolated logical goals, proof engineering entails a broader and more structurally intricate set of tasks, such as extending libraries with new features, restructuring existing proofs (refactoring), and correcting erroneous logic (bug fixing). For a more comprehensive discussion of proof engineering practices and challenges, see Appendix .
Most existing benchmarks emphasise single-goal theorem proving within static contexts, diverging significantly from the demands of real-world proof maintenance. As shown in Table 1, conventional benchmarks evaluate success based on isolated proof completion, neglecting the broader engineering workflows needed to evolve and maintain large formal mathematics libraries. Recent efforts have explored richer proof generation scenarios [4, 6, 11, 12, 5, 13, 1], but still typically focus on isolated reasoning tasks rather than realistic proof engineering workflows. A clear gap remains in systematic, task-diverse benchmarks that reflect the full scope of real-world proof engineering.
@@
Objectives Individual proof completion Feature addition, proof refactoring, bug fixing
Table 1: Comparison between conventional benchmarks and real-world proof engineering practices.
@@ 3 APE-Bench I: A Realistic Benchmark for File-Level Proof Engineering
To concretely realise the paradigm of Automated Proof Engineering (APE), we introduce APE-Bench I, a realistic and instruction-driven benchmark constructed from real-world edits in the Mathlib4 formal mathematics library. This section describes the design and construction of the benchmark, including its task format, the methodology for extracting file-level edits from version history, and the automatic synthesis of natural-language instructions and semantic labels.
@@ 3.1 Task Format and Objective
Each task in APE-Bench I defines a localised file-level proof engineering problem, grounded in an actual edit to a Lean source file. The objective is to apply a valid patch that fulfils a natural-language instruction by transforming a given pre-edit file into a structurally correct and semantically meaningful post-edit version.
@@
  • GoldDiff is a unified diff extracted from human-written commits that serves as the task’s ground truth, specifying changes relative to PreFile. Its format enables modular editing and supports automated application, verification, and evaluation.
This triplet formulation mirrors the practical realities of proof maintenance, wherein developers continuously update and refactor multiple interdependent components within full files, as opposed to resolving isolated subgoals in static settings.
3.2 Task Extraction from Mathlib4 Commits
(a) Token distribution by benchmarks
@@ 3.2 Task Extraction from Mathlib4 Commits
Figure 3: APE-Bench I: Scale and complexity of extracted tasks. (a) APE-Bench I tasks span significantly longer input–output contexts than prior datasets. (b) Starting from 19k commits, the pipeline produces 185k+ candidates, filtered to 10,928 high-quality tasks.
To construct realistic and diverse proof engineering tasks, we extract file-level edits from the version history of Mathlib4. Our extraction pipeline, illustrated in Figure 3(b), proceeds in four stages:
@@
We begin with 19,867 Git commits made between August 2023 and March 2025. Each commit is decomposed into file-level diffs, isolating modified files and their corresponding changes. These diffs are further segmented into semantically scoped chunks by hunk or declaration boundaries.
@@
To improve modelling tractability and enhance semantic interpretability, we decompose large diffs into smaller, contiguous edit fragments. Each fragment is applied incrementally to the evolving file state, yielding fine-grained task units. This expansion increases the total number of edit candidates to over 187,000.
@@
We apply a sequence of automated filters targeting semantic relevance (retaining only commits with recognisable editing intent), content quality (discarding comment-only or whitespace edits), and structural integrity (restricting to at most 100 lines of changes and 16,000 tokens). The overall retention trajectory is visualised in Figure 3(b). For additional details on our filtering methodology, see Appendix B.1.
@@
To ensure syntactic soundness, all retained patches are compiled under their post-edit environments using Lean. Only patches that are successfully compiled in the restored version context are included in the final benchmark.
The final benchmark consists of 10,928 structurally coherent tasks. As shown in Figure 3(a), these tasks exhibit significantly longer input and output sequences than prior benchmarks, reflecting the higher structural complexity of proof engineering tasks.